OpenStudy (anonymous):

I think i did something wrong here -6≤ x/2 -4< -3 how would i write this in interval?

OpenStudy (alexwee123):

well you should simplify it first then try to put it into interval notation :o

OpenStudy (anonymous):

i did i got to different answers :o

OpenStudy (campbell_st):

add 4 to both sides \[-2 \le \frac{x}{2} \le 2\] last step is to multiply each term by 2

OpenStudy (anonymous):

i came up with [-4,2)

OpenStudy (campbell_st):

yes your solution is right I made 2 typos in mine... sorry \[-4\le x < 2\]

OpenStudy (anonymous):

its cool lol and thanks